# Tracing Multilingual Factual Knowledge Acquisition in Pretraining

### v1

@@ 1 Introduction
While prior work has investigated mechanisms of (multilingual) factual recall (Geva et al., 2023; Zhao et al., 2024; Fierro et al., 2024; Liu et al., 2025) and analyzed sources of crosslingual inconsistency (Qi et al., 2023; Wang et al., 2025), these studies have largely focused on final models, drawing conclusions solely from the end of pretraining. As a consequence, the developmental process by which LLMs acquire factual knowledge across languages remains poorly understood.
@@ 2 Related Work
Several studies have investigated how Transformer-based models (Vaswani et al., 2017) acquire linguistic or task-specific knowledge during different phases of pretraining, in both monolingual (Choshen et al., 2022; Xia et al., 2023; Müller-Eberstein et al., 2023; Chen et al., 2024) and multilingual settings (Blevins et al., 2022; Wang et al., 2024). A concurrent study by Merullo et al. (2025) most closely resembles our work; they demonstrate that fact frequency is a strong predictor of both factual recall and the emergence of linear factual representations (e.g., subject-to-object mappings via linear transformation) (Hernandez et al., 2024). However, their analysis is conducted in a purely monolingual context. In contrast, our work examines multilingual factual knowledge acquisition and shows that while fact frequency remains a key driver of factual recall, crosslingual knowledge transfer provides additional – albeit limited – benefits in enhancing multilingual factual recall.
@@ 3.2 Multilingual Factual Dataset
We use KLAR (Wang et al., 2025), a multilingual factual knowledge probing dataset, for our investigation. We use 1,197 facts grouped into 12 relation categories (cf. Table 2 in §A). Each fact is represented as a triple of subject, relation, and object. KLAR also provides a prompt template for each relation in each language, structured as “<Question> The answer is:”. For example, for triple (France, capital, Paris), the template will then be expanded as “Where is France’s capital located? The answer is:”, with expected answer “Paris” in English. All facts and prompt templates are available in all 12 languages. We therefore transform each fact into a query $q_{i}^{l}$ with expected answer $o_{i}^{l}$ in language $l$; for each fact $i$, $q_{i}^{l}$ and $q_{i}^{l^{\prime}}$ are translations of the same query in languages $l$ and $l^{\prime}$. We denote the resulting set of queries as $Q$.
@@ 3.4 Fact Frequencies
We approximate a fact’s frequency by counting the number of documents where its subject and object co-occur in the pretraining corpus. This co-occurrence-based approximation has been widely used and shown to be reliable (Elsahar et al., 2018; Elazar et al., 2023; Merullo et al., 2025; Liu et al., 2025). For some languages, this approximation is fairly accurate due to the uniqueness of their scripts – for example, the subject-object pair ({CJK}UTF8gbsn法国, {CJK}UTF8gbsn巴黎) in Chinese is unlikely to appear in texts from other languages. However, ambiguity arises in languages that share scripts, such as English and French. The same pair (France, Paris), for instance, may appear in either language, resulting in an aggregated frequency count shared across both. We analyze the impact of this identical-fact effect and show that it does not compromise the robustness of our findings (cf. §I). To efficiently obtain these co-occurrence counts, we use the ElasticSearch API provided by WIMBD (Elazar et al., 2024), a tool designed for scalable search and frequency analysis over large corpora. [^4] A public demo of WIMBD is available at: https://wimbd.apps.allenai.org/. All fact frequencies in our analysis are computed over the Dolma v1.7 corpus (Soldaini et al., 2024) used to pretrain OLMo, by measuring the number of subject-object co-occurrences for each fact in KLAR.
@@ 4 Multilingual Factual Recall Dynamics
[…]
@@ 5.3 Recall Prediction with Frequencies
Formally, for each language $l$, we define a dataset $\mathcal{D}_{l}=\{(f_{i}^{l},y_{i}^{l})\}_{i=1}^{N}$, where $f_{i}^{l}\in\mathbb{Z}_{\geq 0}$ is the frequency of fact $i$, and $y_{i}^{l}\in\{0,1\}$ indicates whether the model correctly recalled the fact ($1$ if correct, 00 otherwise). $\mathbb{Z}_{\geq 0}$ is the set of positive integers including 0. We then define a threshold-based classifier $h_{t}^{l}(f)$ for each language as: $h_{t}^{l}(f)=\begin{cases}1,&\text{if }f\geq t\\ 0,&\text{otherwise}\end{cases}$. The optimal threshold $t^{*}_{l}$ in each language is selected to maximize classification accuracy:
@@
where $\mathbf{1}(\cdot)$ is the indicator function. To better understand the classification behavior, we also compute the number of false negatives (FN) under the optimal threshold, as these facts are also correctly predicted but with low frequencies.555 Other error types are not the primary focus of our further analysis presented in the main content. For example, the cause of false positives may be due to (1) insufficient exposure to the fact despite its high frequency, or (2) sensitivity to the specific prompt used for evaluation. We present an analysis of the classifier in §E and complete error breakdown in §I. Table 1 presents the classification performance.
@@ 6.3 Similarity Dynamics
To better understand why certain languages, particularly those that do not use the Latin script, benefit from knowledge acquired in English, we analyze the evolution of cosine similarity between sentence-level representations of prompts (cf. §3.2) or fact pairs corresponding to SCLFPs during pretraining. Specifically, we create fact pairs of SCLFPs for each language, where every pair contains one prompt in that language and its counterpart in English. We then track the cosine similarity between these paired representations across checkpoints. [^7] We use the contextualized embedding of the final token as the sentence-level representation. Representations are extracted at each layer, and we report the mean cosine similarity computed by averaging similarities across all layers. As a baseline, we also compute cosine similarities for UWLFPs – unsurprisingly wrong low-frequency predictions identified in our frequency-based classification (cf. §5.2) – as well as for all fact pairs in each language. Figure 7 illustrates the progression of similarity scores over time for 6 languages (full results are available in §F).888 To avoid inflated similarity, for each language, we filter out fact pairs where the object strings in that language and English are identical. Table 4 in §H provides statistics of fact pairs containing identical objects across languages.
@@ Appendix E Threshold Classifier Sensitivity
Figure 13: Classifier accuracy versus selected frequency threshold within a range of ±$\pm 20\%$ of $t_{l}^{*}$, the chosen threshold. The dotted line shows the actual chosen threshold.
@@ H.1 Same Object Effect
Figure 16: Distribution of identical-object facts across relation types for each language. A cell labeled “$17/24$” indicates that 17 out of 24 facts are correctly recalled, where the 24 facts are those whose English counterparts are also correctly predicted. Cells marked “$/0$ 0” indicate that no such facts exist for that relation in the given language. The results suggest that many languages, particularly those using the Latin script, benefit from sharing identical object strings with English.
@@ Appendix I Effects of Excluding Identical Facts Across Languages
Appendix L Experimental Environment and Hyperparameters

### v2

@@ 1 Introduction
While prior work has investigated mechanisms of (multilingual) factual recall (Geva et al., 2023; Zhao et al., 2024; Chang et al., 2024; Fierro et al., 2024; Liu et al., 2025b) and analyzed sources of crosslingual inconsistency (Qi et al., 2023; Wang et al., 2025), these studies have largely focused on final models, drawing conclusions solely from the end of pretraining. As a result, the developmental process by which LLMs acquire factual knowledge across languages remains poorly understood.
@@ 2 Related Work
Several studies have investigated how Transformer-based models (Vaswani et al., 2017) acquire linguistic or task-specific knowledge during different phases of pretraining, in both monolingual (Choshen et al., 2022; Xia et al., 2023; Müller-Eberstein et al., 2023; Chen et al., 2024; Chang et al., 2024) and multilingual settings (Blevins et al., 2022; Wang et al., 2024). A concurrent study by Merullo et al. (2025) most closely resembles our work; they demonstrate that fact frequency is a strong predictor of both factual recall and the emergence of linear factual representations (e.g., subject-to-object mappings via linear transformation) (Hernandez et al., 2024). However, their analysis is conducted in a purely monolingual context. In contrast, our work examines multilingual factual knowledge acquisition and shows that while fact frequency remains a key driver of factual recall, crosslingual knowledge transfer provides additional – albeit limited – benefits in enhancing multilingual factual recall.
@@ 3.2 Multilingual Factual Dataset
We use KLAR (Wang et al., 2025), a multilingual factual knowledge probing dataset, for our investigation. We use 1,197 facts grouped into 12 relation categories (cf. Table 2 in §A). Each fact is represented as a triple of subject, relation, and object. KLAR also provides a prompt template for each relation in each language, structured as “<Question> The answer is:”. For example, for triple (France, capital, Paris), the template will then be expanded as “Where is France’s capital located? The answer is:”, with expected answer “Paris” in English. All facts and prompt templates are available in all 12 languages. We therefore transform each fact into a query $q_{i}^{l}$ with expected answer $o_{i}^{l}$ in language $l$; for each fact $i$, $q_{i}^{l}$ and $q_{i}^{l^{\prime}}$ are translations of the same query in languages $l$ and $l^{\prime}$. We denote the resulting set of queries as QQ. [^2] In this paper, we only consider a single prompt template for each relation in each language, since the influence of prompt variation is expected to be limited due to the simplicity of factual recall. We present an additional study on how different prompt templates affect the factual recall in §L.
@@ 3.4 Fact Frequencies
We approximate a fact’s frequency by counting the number of documents where its subject and object co-occur in the pretraining corpus. This co-occurrence-based approximation has been widely used and shown to be reliable (Elsahar et al., 2018; Elazar et al., 2023; Merullo et al., 2025; Liu et al., 2025b). For some languages, this approximation is fairly accurate due to the uniqueness of their scripts – for example, the subject-object pair (法国, 巴黎) in Chinese is unlikely to appear in texts from other languages. However, ambiguity arises in languages that share scripts, such as English and French. The same pair (France, Paris), for instance, may appear in either language, resulting in an aggregated frequency count shared across both. We analyze the impact of this identical-fact effect and show that it does not compromise the robustness of our findings (cf. §I). To efficiently obtain these co-occurrence counts, we use the ElasticSearch API provided by WIMBD (Elazar et al., 2024), a tool designed for scalable search and frequency analysis over large corpora. [^5] A public demo of WIMBD is available at: https://wimbd.apps.allenai.org/. An alternative with similar functionality is Infini-gram (Liu et al., 2025a): https://infini-gram.readthedocs.io/en/latest/api.html. All fact frequencies in our analysis are computed over the Dolma v1.7 corpus (Soldaini et al., 2024) used to pretrain OLMo, by measuring the number of subject-object co-occurrences for each fact in KLAR.
@@ 4 Multilingual Factual Recall Dynamics
We observe that the trajectory of crosslingual consistency in each language $l\neq\text{eng\_Latn}$ closely mirrors its own factual accuracy throughout pretraining. This suggests that consistency is primarily driven by whether the fact is correctly recalled in $l$, which almost always implies that it is also recalled in English. The implication is twofold. (1) For non-English languages, the consistency of a language (CO) is effectively gated by its performance (ACC). (2) The limited capability of the model to transfer knowledge from English to other languages, referred to as the crosslingual knowledge barrier (Chua et al., 2025), is a persistent problem throughout pretraining.
@@ 5.3 Recall Prediction with Frequencies
Formally, for each language $l$, we define a dataset $\mathcal{D}_{l}=\{(f_{i}^{l},y_{i}^{l})\}_{i=1}^{N}$, where $f_{i}^{l}\in\mathbb{Z}_{\geq 0}$ is the frequency of fact $i$, and $y_{i}^{l}\in\{0,1\}$ indicates whether the model correctly recalled the fact ($1$ if correct, 0 otherwise). $\mathbb{Z}_{\geq 0}$ is the set of positive integers including 0. We then define a threshold-based classifier $h_{t}^{l}(f)$ for each language as: $h_{t}^{l}(f)=\begin{cases}1,&\text{if }f\geq t\\ 0,&\text{otherwise}\end{cases}$. The optimal threshold $t^{*}_{l}$ in each language is selected to maximize classification accuracy:
@@
where $\mathbf{1}(\cdot)$ is the indicator function. To better understand the classification behavior, we also compute the number of false negatives (FN) under the optimal threshold, as these facts are also correctly predicted but with low frequencies. [^6] Other error types are not the primary focus of our further analysis presented in the main content. For example, the cause of false positives may be due to (1) insufficient exposure to the fact despite its high frequency, or (2) sensitivity to the specific prompt used for evaluation. We present an analysis of the classifier in §E and a complete error breakdown in §I. Table 1 presents the classification performance.
@@ 6.3 Similarity Dynamics
To better understand why certain languages, particularly those that do not use the Latin script, benefit from knowledge acquired in English, we analyze the evolution of cosine similarity between sentence-level representations of prompts (cf. §3.2) or fact pairs corresponding to SCLFPs during pretraining. Specifically, we create fact pairs of SCLFPs for each language, where every pair contains one prompt in that language and its counterpart in English. We then track the cosine similarity between these paired representations across checkpoints. [^8] We use the contextualized embedding of the final token as the sentence-level representation. Representations are extracted at each layer, and we report the mean cosine similarity computed by averaging similarities across all layers. As a baseline, we also compute cosine similarities for UWLFPs – unsurprisingly wrong low-frequency predictions identified in our frequency-based classification (cf. §5.2) – as well as for all fact pairs in each language. Figure 7 illustrates the progression of similarity scores over time for 6 languages (full results are available in §F). [^9] To avoid inflated similarity, for each language, we filter out fact pairs where the object strings in that language and English are identical. Table 4 in §H provides statistics of fact pairs containing identical objects across languages.
@@ Appendix E Threshold Classifier Sensitivity
Figure 13: Classifier accuracy versus selected frequency threshold within a range of $\pm 20\%$ of $t_{l}^{*}$, the chosen threshold. The dotted line shows the actual chosen threshold.
@@ H.1 Same Object Effect
Figure 16: Distribution of identical-object facts across relation types for each language. A cell labeled “17/2417/24” indicates that 17 out of 24 facts are correctly recalled, where the 24 facts are those whose English counterparts are also correctly predicted. Cells marked “0/00/0” indicate that no such facts exist for that relation in the given language. The results suggest that many languages, particularly those using the Latin script, benefit from sharing identical object strings with English.
@@ Appendix I Effects of Excluding Identical Facts Across Languages
Factual Recall (ACC, %) Consistency (CO, %)
Language Template 1 Template 2 Template 3 Template 4 Template 5 Template 1 Template 2 Template 3 Template 4 Template 5
ara_Arab 17.5 18.2 18.1 18.3 18.1 18.9 19.3 19.5 19.5 18.9
cat_Latn 44.2 45.3 43.5 45.4 44.7 48.5 49.4 47.6 49.3 48.5
ell_Grek 18.8 17.1 18.6 17.2 18.8 20.0 18.0 19.7 18.2 19.7
eng_Latn 82.0 81.9 81.4 81.9 82.2 100.0 100.0 100.0 100.0 100.0
spa_Latn 56.4 55.6 55.4 55.7 56.8 62.0 60.2 61.8 61.2 61.7
fra_Latn 62.0 62.7 62.7 62.0 61.3 68.1 68.1 69.3 67.8 67.5
jpn_Jpan 18.5 18.1 17.8 19.5 20.3 19.6 19.1 18.3 20.6 21.2
kor_Kore 20.0 18.9 18.8 19.8 19.1 20.7 19.3 19.4 20.8 20.5
rus_Cyrl 31.2 31.1 32.1 30.6 32.2 33.3 32.6 33.9 31.7 33.5
tur_Latn 40.1 40.8 40.4 41.3 39.9 42.3 43.4 42.8 43.8 42.7
ukr_Cyrl 23.7 23.1 23.3 23.1 24.0 24.8 24.1 24.2 23.9 25.2
zho_Hans 26.2 25.2 26.3 26.2 25.1 27.6 26.7 27.5 27.1 26.1
Table 6: Factual recall accuracy and crosslingual consistency (with respect to English) across five prompt templates.
Appendix L Effect of Prompt Template Variation
Prompt template variation, or prompt phrasing, can largely affect LLM behavior, particularly for open-ended or generative tasks. However, in the context of factual recall, where the expected output is typically a short, well-defined answer, the influence of prompt variation is more limited. To verify this, we conduct a case study using 5 different prompt templates provided by KLAR (Wang et al., 2025) (Template 1 in them is used in the main text of this paper) on the step-400000 checkpoint. The results of factual recall accuracy and consistency (with respect to English) are shown in Table 6. We can observe that the factual recall performance (accuracy and consistency) remains consistent across different prompts in all languages, confirming that the prompt variation does not have a substantial effect on the factual recall. These complementary results indicate the robustness of our findings in the main text.
Appendix M Experimental Environment and Hyperparameters